# The first phase of mass transfer in low-mass binaries: neither stable nor a common envelope

### v1

@@ Acknowledgements.
- Amaro-Seoane et al. (2023) Amaro-Seoane, P., Andrews, J., Arca Sedda, M., et al. 2023, Living Reviews in Relativity, 26, 2
- Bonačić Marinović et al. (2008) Bonačić Marinović, A. A., Glebbeek, E., & Pols, O. R. 2008, A&A, 480, 797
- Bours et al. (2014) Bours, M. C. P., Marsh, T. R., Parsons, S. G., et al. 2014, MNRAS, 438, 3399
- Brown et al. (2020) Brown, W. R., Kilic, M., Bédard, A., Kosakowski, A., & Bergeron, P. 2020, ApJ, 892, L35
- Burdge et al. (2020a) Burdge, K. B., Coughlin, M. W., Fuller, J., et al. 2020a, ApJ, 905, L7
- Burdge et al. (2020b) Burdge, K. B., Prince, T. A., Fuller, J., et al. 2020b, ApJ, 905, 32
- Chen et al. (2022) Chen, H.-L., Tauris, T. M., Chen, X., & Han, Z. 2022, The Astrophysical Journal, 925, 89
- Chen et al. (2017) Chen, X., Maxted, P. F. L., Li, J., & Han, Z. 2017, MNRAS, 467, 1874
- De Marco et al. (2011) De Marco, O., Passy, J.-C., Moe, M., et al. 2011, MNRAS, 411, 2277
- Dermine et al. (2013) Dermine, T., Izzard, R. G., Jorissen, A., & Van Winckel, H. 2013, A&A, 551, A50
- Dewberry & Wu (2025) Dewberry, J. W. & Wu, Y. 2025, ApJ, 984, 137
- Di Stefano et al. (2023) Di Stefano, R., Kruckow, M. U., Gao, Y., Neunteufel, P. G., & Kobayashi, C. 2023, ApJ, 944, 87
- Eggleton (1983) Eggleton, P. P. 1983, ApJ, 268, 368
- Escorza et al. (2020) Escorza, A., Siess, L., Van Winckel, H., & Jorissen, A. 2020, A&A, 639, A24
- Gao & Li (2023) Gao, S.-J. & Li, X.-D. 2023, MNRAS, 525, 2605
- Ge et al. (2024) Ge, H., Tout, C. A., Webbink, R. F., et al. 2024, ApJ, 961, 202
- Ge et al. (2020) Ge, H., Webbink, R. F., Chen, X., & Han, Z. 2020, ApJ, 899, 132
- Hallakoun et al. (2016) Hallakoun, N., Maoz, D., Kilic, M., et al. 2016, MNRAS, 458, 845
- Han (1998) Han, Z. 1998, MNRAS, 296, 1019
- Han et al. (1995) Han, Z., Podsiadlowski, P., & Eggleton, P. P. 1995, MNRAS, 272, 800
- Iben et al. (1997) Iben, Jr., I., Tutukov, A. V., & Yungelson, L. R. 1997, ApJ, 475, 291
- Jorissen (2019) Jorissen, A. 2019, Mem. Soc. Astron. Italiana, 90, 395
- Kilic et al. (2020) Kilic, M., Bédard, A., Bergeron, P., & Kosakowski, A. 2020, MNRAS, 493, 2805
- Kilic et al. (2021) Kilic, M., Brown, W. R., Bédard, A., & Kosakowski, A. 2021, ApJ, 918, L14
- Korol et al. (2017) Korol, V., Rossi, E. M., Groot, P. J., et al. 2017, MNRAS, 470, 1894
- Kosakowski et al. (2023) Kosakowski, A., Brown, W. R., Kilic, M., et al. 2023, ApJ, 950, 141
- Lamberts et al. (2019) Lamberts, A., Blunt, S., Littenberg, T. B., et al. 2019, MNRAS, 490, 5888
- Li et al. (2023) Li, Z., Chen, X., Ge, H., Chen, H.-L., & Han, Z. 2023, A&A, 669, A82
- Lin et al. (2011) Lin, J., Rappaport, S., Podsiadlowski, P., et al. 2011, ApJ, 732, 70
- Liu et al. (2010) Liu, J., Zhang, Y., Han, Z., & Zhang, F. 2010, Ap&SS, 329, 297
- Liu et al. (2023) Liu, Z.-W., Röpke, F. K., & Han, Z. 2023, Research in Astronomy and Astrophysics, 23, 082001
- Maoz et al. (2014) Maoz, D., Mannucci, F., & Nelemans, G. 2014, ARA&A, 52, 107
- Marsh (1995) Marsh, T. R. 1995, MNRAS, 275, L1
- Maxted et al. (2002) Maxted, P. F. L., Marsh, T. R., & Moran, C. K. J. 2002, MNRAS, 332, 745
- Moran et al. (1997) Moran, C., Marsh, T. R., & Bragaglia, A. 1997, MNRAS, 288, 538
@@
- Moran et al. (1999) Moran, C., Marsh, T. R., & Maxted, P. F. L. 1999, in Astronomical Society of the Pacific Conference Series, Vol. 169, 11th European Workshop on White Dwarfs, ed. S. E. Solheim & E. G. Meistas, 275
- Munday et al. (2024) Munday, J., Pelisoli, I., Tremblay, P. E., et al. 2024, MNRAS, 532, 2534
- Munday et al. (2025) Munday, J., Pelisoli, I., Tremblay, P. E., et al. 2025, MNRAS, 540, 1272
- Munday et al. (2023) Munday, J., Tremblay, P. E., Hermes, J. J., et al. 2023, MNRAS, 525, 1814
- Napiwotzki et al. (2020) Napiwotzki, R., Karl, C. A., Lisker, T., et al. 2020, A&A, 638, A131
- Nelemans et al. (2004) Nelemans, G., Jonker, P. G., Marsh, T. R., & van der Klis, M. 2004, MNRAS, 348, L7
- Nelemans et al. (2000) Nelemans, G., Verbunt, F., Yungelson, L. R., & Portegies Zwart, S. F. 2000, A&A, 360, 1011
- Nelemans et al. (2001) Nelemans, G., Yungelson, L. R., Portegies Zwart, S. F., & Verbunt, F. 2001, A&A, 365, 491
- Oomen et al. (2020) Oomen, G.-M., Pols, O., Van Winckel, H., & Nelemans, G. 2020, A&A, 642, A234
- Paczyński (1971) Paczyński, B. 1971, ARA&A, 9, 183
@@
- Paczynski (1976) Paczynski, B. 1976, in Structure and Evolution of Close Binary Systems, ed. P. Eggleton, S. Mitton, & J. Whelan, Vol. 73, 75
- Panei et al. (2007) Panei, J. A., Althaus, L. G., Chen, X., & Han, Z. 2007, MNRAS, 382, 779
- Parsons et al. (2020) Parsons, S. G., Brown, A. J., Littlefair, S. P., et al. 2020, Nature Astronomy, 4, 690
- Parsons et al. (2023) Parsons, S. G., Hernandez, M. S., Toloza, O., et al. 2023, MNRAS, 518, 4579
- Rappaport et al. (1995) Rappaport, S., Podsiadlowski, P., Joss, P. C., Di Stefano, R., & Han, Z. 1995, MNRAS, 273, 731
- Refsdal & Weigert (1971) Refsdal, S. & Weigert, A. 1971, A&A, 13, 367
- Ruiter et al. (2010) Ruiter, A. J., Belczynski, K., Benacquista, M., Larson, S. L., & Williams, G. 2010, ApJ, 717, 1006
- Ruiter & Seitenzahl (2024) Ruiter, A. J. & Seitenzahl, I. R. 2024, arXiv e-prints, arXiv:2412.01766
- Saladino & Pols (2019) Saladino, M. I. & Pols, O. R. 2019, A&A, 629, A103
- Schaffenroth et al. (2023) Schaffenroth, V., Barlow, B. N., Pelisoli, I., Geier, S., & Kupfer, T. 2023, A&A, 673, A90
- Schaffenroth et al. (2022) Schaffenroth, V., Pelisoli, I., Barlow, B. N., Geier, S., & Kupfer, T. 2022, A&A, 666, A182
- Scherbak & Fuller (2023) Scherbak, P. & Fuller, J. 2023, MNRAS, 518, 3966
- Snaith et al. (2015) Snaith, O., Haywood, M., Di Matteo, P., et al. 2015, A&A, 578, A87
- Stoppa et al. (2023) Stoppa, F., Cator, E., & Nelemans, G. 2023, MNRAS, 524, 1061
- Tang et al. (2024) Tang, P., Eldridge, J. J., Meyer, R., et al. 2024, MNRAS, 534, 1707
- Tauris & Savonije (1999) Tauris, T. M. & Savonije, G. J. 1999, A&A, 350, 928
- Temmink et al. (2023) Temmink, K. D., Pols, O. R., Justham, S., Istrate, A. G., & Toonen, S. 2023, A&A, 669, A45
- Toonen et al. (2012) Toonen, S., Nelemans, G., & Portegies Zwart, S. 2012, A&A, 546, A70
- Tutukov & Yungelson (1988) Tutukov, A. V. & Yungelson, L. R. 1988, Soviet Astronomy Letters, 14, 265
- Vos et al. (2015) Vos, J., Østensen, R. H., Marchant, P., & Van Winckel, H. 2015, A&A, 579, A49
@@
- Webbink (2008) Webbink, R. F. 2008, in Astrophysics and Space Science Library, Vol. 352, Astrophysics and Space Science Library, ed. E. F. Milone, D. A. Leahy, & D. W. Hobill, 233
- Wesemael et al. (1994) Wesemael, F., Bergeron, P., Lamontagne, R. L., et al. 1994, ApJ, 429, 369
- Woods et al. (2012) Woods, T. E., Ivanova, N., van der Sluys, M. V., & Chaichenets, S. 2012, ApJ, 744, 12
- Yu & Jeffery (2010) Yu, S. & Jeffery, C. S. 2010, A&A, 521, A85
- Zhang et al. (2021) Zhang, Y., Chen, H.-L., Xiong, H., Chen, X., & Han, Z. 2021, MNRAS, 505, 3514
- Zhang et al. (2024) Zhang, Y., Li, Z., Chen, X., & Han, Z. 2024, ApJ, 977, 24
@@ Appendix A Period/radius - core mass relations
Several studies, in particular into the formation of millisecond radio pulsars with WD companions have determined the period – core mass/WD mass relation that is expected after a phase of stable mass transfer. Rappaport et al. (1995) gave a simple relation, but Tauris & Savonije (1999) and Lin et al. (2011) found that at low mass that relation overestimates the periods. In addition, the radius of the giants and thus the period – core mass relation depends on metallicity. Rappaport et al. (1995) and Tauris & Savonije (1999) find that population II stars, with assumed metallicity of 0.001, i.e about 1/20 of Solar, have about 40% smaller radii. However, such low metallicity is very rare in the Milky Way. In Fig. 7 we plot the cumulative distribution of the metallicity of the total amount of star formation in the Milky Way, based on the derived star formation rate and metallicity by Snaith et al. (2015). It can be seen that there is a similar amount of star formation with metallicities higher than Solar as there is with metallicities below , which on a logarithmic scale lies between the two metallicities used in the models by Tauris & Savonije (1999). We therefore use two metallicities in this study: Solar and intermediate in which the radii are 20% smaller than the Solar values. The latter is similar to ”Pop I+II” in Tauris & Savonije (1999), For the population studies we also use these two metallicities as metallicity bins. The lower metallicity covers about 1/3 of the total SFR (see Fig. 7).
We adapted the curve of Lin et al. (2011) that is given only for solar metallicity and for relatively low mass to match the ones of Rappaport et al. (1995); Tauris & Savonije (1999) at higher mass. For this we multiply their relation with a factor 1.1 for solar metallicity and a factor 0.8 for intermediate metallicity. We note that Gao & Li (2023) find somewhat larger periods.

### v2

@@ Acknowledgements.
- Astrophysics with the Laser Interferometer Space Antenna. Living Reviews in Relativity 26 (1), pp. 2. External Links: Document, 2203.06016, ADS entry Cited by: §1.
- Orbital eccentricities of binary systems with a former AGB star. A&A 480 (3), pp. 797–805. External Links: Document, 0710.4859, ADS entry Cited by: §5.
- Precise parameters for both white dwarfs in the eclipsing binary CSS 41177. MNRAS 438 (4), pp. 3399–3408. External Links: Document, 1401.1503, ADS entry Cited by: Table 1.
- A 1201 s Orbital Period Detached Binary: The First Double Helium Core White Dwarf LISA Verification Binary. ApJ 892 (2), pp. L35. External Links: Document, 2004.00641, ADS entry Cited by: Table 1.
- An 8.8 Minute Orbital Period Eclipsing Detached Double White Dwarf Binary. ApJ 905 (1), pp. L7. External Links: Document, 2010.03555, ADS entry Cited by: Table 1.
- A Systematic Search of Zwicky Transient Facility Data for Ultracompact Binary LISA-detectable Gravitational-wave Sources. ApJ 905 (1), pp. 32. External Links: Document, 2009.02567, ADS entry Cited by: Table 1, Table 1, Table 1, Table 1, Table 1, Table 1.
- Formation of the double white dwarf binary ptf j0533+0209 through stable mass transfer?. The Astrophysical Journal 925 (1), pp. 89. External Links: ISSN 0004-637X, 1538-4357, Document Cited by: §2.
- The Formation of EL CVn-type Binaries. MNRAS 467 (2), pp. 1874–1889. External Links: Document, 1604.01956, ADS entry Cited by: §2.
- On the formalism for the common envelope interaction. MNRAS 411 (4), pp. 2277–2292. External Links: Document, 1010.4374, ADS entry Cited by: §1.
- Eccentricity-pumping in post-AGB stars with circumbinary discs. A&A 551, pp. A50. External Links: Document, ADS entry Cited by: §5.
- Testing Tidal Theory Using Gaia Binaries: The Red Giant Branch. ApJ 984 (2), pp. 137. External Links: Document, 2501.13929, ADS entry Cited by: §5.
- SCATTER: A New Common Envelope Formalism. ApJ 944 (1), pp. 87. External Links: Document, 2212.06770, ADS entry Cited by: Figure 6, §5.
- Aproximations to the radii of Roche lobes.. ApJ 268, pp. 368–369. External Links: Document, ADS entry Cited by: §2.1.
- Binary evolution along the red giant branch with BINSTAR: The barium star perspective. A&A 639, pp. A24. External Links: Document, 2005.05391, ADS entry Cited by: §5.
- The white dwarf mass-orbital period relation under wind mass-loss. MNRAS 525 (2), pp. 2605–2615. External Links: Document, 2308.05641, ADS entry Cited by: Appendix A, §5.
- The Common Envelope Evolution Outcome. II. Short-orbital-period Hot Subdwarf B Binaries Reveal a Clear Picture. ApJ 961 (2), pp. 202. External Links: Document, 2311.17304, ADS entry Cited by: §5.
- Adiabatic Mass Loss in Binary Stars. III. From the Base of the Red Giant Branch to the Tip of the Asymptotic Giant Branch. ApJ 899 (2), pp. 132. External Links: Document, 2007.09848, ADS entry Cited by: §1.
- SDSS J1152+0248: an eclipsing double white dwarf from the Kepler K2 campaign. MNRAS 458 (1), pp. 845–854. External Links: Document, 1507.06311, ADS entry Cited by: Table 1.
- The formation of bipolar planetary nebulae and close white dwarf binaries. MNRAS 272 (4), pp. 800–820. External Links: Document, ADS entry Cited by: §1.
- The formation of double degenerates and related objects. MNRAS 296 (4), pp. 1019–1040. External Links: Document, ADS entry Cited by: §1.
- Helium and Carbon-Oxygen White Dwarfs in Close Binaries. ApJ 475 (1), pp. 291–299. External Links: Document, ADS entry Cited by: §1.
- Impact of binaries on stellar evolution in the Gaia era. Mem. Soc. Astron. Italiana 90, pp. 395. External Links: ADS entry Cited by: §1.
- Two new double-lined spectroscopic binary white dwarfs. MNRAS 493 (2), pp. 2805–2816. External Links: Document, 2002.06214, ADS entry Cited by: Table 1.
- The Discovery of Two LISA Sources within 0.5 kpc. ApJ 918 (1), pp. L14. External Links: Document, 2108.08324, ADS entry Cited by: Table 1, Table 1.
- Prospects for detection of detached double white dwarf binaries with Gaia, LSST and LISA. MNRAS 470 (2), pp. 1894–1910. External Links: Document, 1703.02555, ADS entry Cited by: §1.
- The ELM Survey South. II. Two Dozen New Low-mass White Dwarf Binaries. ApJ 950 (2), pp. 141. External Links: Document, 2305.03079, ADS entry Cited by: Table 1.
- Predicting the LISA white dwarf binary population in the Milky Way with cosmological simulations. MNRAS 490 (4), pp. 5888–5903. External Links: Document, 1907.00014, ADS entry Cited by: §1.
- Influence of a mass transfer stability criterion on double white dwarf populations. A&A 669, pp. A82. External Links: Document, 2211.01861, ADS entry Cited by: §1, §5.
- LMXB and IMXB Evolution: I. The Binary Radio Pulsar PSR J1614-2230. ApJ 732 (2), pp. 70. External Links: Document, 1012.1877, ADS entry Cited by: Figure 8, Appendix A, Appendix A, Appendix A, §2.1, §5.
- Contribution of close double white dwarfs to the Galactic gravitational wave foreground. Ap&SS 329 (1-2), pp. 297–303. External Links: Document, ADS entry Cited by: §1.
- Type Ia Supernova Explosions in Binary Systems: A Review. Research in Astronomy and Astrophysics 23 (8), pp. 082001. External Links: Document, 2305.13305, ADS entry Cited by: §1.
- Observational Clues to the Progenitors of Type Ia Supernovae. ARA&A 52, pp. 107–170. External Links: Document, 1312.0628, ADS entry Cited by: §1.
- The discovery of a short-period double-degenerate binary star. MNRAS 275 (1), pp. L1–L5. External Links: Document, astro-ph/9504088, ADS entry Cited by: Table 1.
- The mass ratio distribution of short-period double degenerate stars. MNRAS 332 (3), pp. 745–753. External Links: Document, astro-ph/0201411, ADS entry Cited by: Table 1, Table 1.
- A detached double degenerate with a 1.4-h orbital period. MNRAS 288 (2), pp. 538–544. External Links: Document, astro-ph/9704240, ADS entry Cited by: Table 1.
@@
- Mass ratios of double degenerate binaries. In 11th European Workshop on White Dwarfs, S. -E. Solheim and E. G. Meistas (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 169, pp. 275. External Links: ADS entry Cited by: §1.
- The DBL Survey I: discovery of 34 double-lined double white dwarf binaries. MNRAS 532 (2), pp. 2534–2556. External Links: Document, 2407.02594, ADS entry Cited by: Table 1, §3.
- Correction to: The DBL Survey I: discovery of 34 double-lined double white dwarf binaries. MNRAS 540 (1), pp. 1272–1276. External Links: Document, ADS entry Cited by: Table 1, Table 1, Table 1, Table 1, Table 1, Table 1, Table 1, Table 1, Table 1, Table 1, Table 1, §3.
- An eclipsing 47 min double white dwarf binary at 400 pc. MNRAS 525 (2), pp. 1814–1823. External Links: Document, 2308.00036, ADS entry Cited by: Table 1.
- The ESO supernovae type Ia progenitor survey (SPY). The radial velocities of 643 DA white dwarfs. A&A 638, pp. A131. External Links: Document, 1906.10977, ADS entry Cited by: Table 1.
- Reconstructing the evolution of white dwarf binaries: further evidence for an alternative algorithm for the outcome of the common-envelope phase in close binaries. MNRAS 356 (2), pp. 753–764. External Links: Document, astro-ph/0410301, ADS entry Cited by: §2.2, §2.3.
- Reconstructing the evolution of double helium white dwarfs: envelope loss without spiral-in. A&A 360, pp. 1011–1018. External Links: Document, astro-ph/0006216, ADS entry Cited by: §1, §1, §1, §2.1, §2.3.
- Population synthesis for double white dwarfs . I. Close detached systems. A&A 365, pp. 491–507. External Links: Document, astro-ph/0010457, ADS entry Cited by: §1, footnote 1.
- Disc-binary interactions in depleted post-AGB binaries. A&A 642, pp. A234. External Links: Document, 2008.08097, ADS entry Cited by: §5.
- Evolutionary Processes in Close Binary Systems. ARA&A 9, pp. 183. External Links: Document, ADS entry Cited by: §1.
@@
- Common Envelope Binaries. In Structure and Evolution of Close Binary Systems, P. Eggleton, S. Mitton, and J. Whelan (Eds.), Vol. 73, pp. 75. External Links: ADS entry Cited by: §1.
- Full evolution of low-mass white dwarfs with helium and oxygen cores. MNRAS 382 (2), pp. 779–792. External Links: Document, ADS entry Cited by: §3.
- The white dwarf binary pathways survey - IX. Three long period white dwarf plus subgiant binaries. MNRAS 518 (3), pp. 4579–4594. External Links: Document, 2211.08440, ADS entry Cited by: §1.
- A pulsating white dwarf in an eclipsing binary. Nature Astronomy 4, pp. 690–696. External Links: Document, 2003.07371, ADS entry Cited by: Table 1.
- The relation between white dwarf mass and orbital period in wide binary radio pulsars. MNRAS 273 (3), pp. 731–741. External Links: Document, ADS entry Cited by: Figure 8, Appendix A, Appendix A, Appendix A, §2.1, §5.
- On the Production of White Dwarfs in Binary Systems of Small Mass. A&A 13, pp. 367. External Links: ADS entry Cited by: §1.
- The LISA Gravitational Wave Foreground: A Study of Double White Dwarfs. ApJ 717 (2), pp. 1006–1021. External Links: Document, 0705.3272, ADS entry Cited by: §1.
- Type Ia supernova progenitors: a contemporary view of a long-standing puzzle. arXiv e-prints, pp. arXiv:2412.01766. External Links: Document, 2412.01766, ADS entry Cited by: §1.
- The eccentric behaviour of windy binary stars. A&A 629, pp. A103. External Links: Document, 1906.02038, ADS entry Cited by: §5.
- Hot subdwarfs in close binaries observed from space. II. Analyses of the light variations. A&A 673, pp. A90. External Links: Document, 2302.12507, ADS entry Cited by: §5.
- Hot subdwarfs in close binaries observed from space. I. Orbital, atmospheric, and absolute parameters, and the nature of their companions. A&A 666, pp. A182. External Links: Document, 2207.02001, ADS entry Cited by: §5.
- White dwarf binaries suggest a common envelope efficiency 1/3. MNRAS 518 (3), pp. 3966–3984. External Links: Document, 2211.02036, ADS entry Cited by: §1.
- Reconstructing the star formation history of the Milky Way disc(s) from chemical abundances. A&A 578, pp. A87. External Links: Document, 1410.3829, ADS entry Cited by: Figure 7, Appendix A.
- Consistency tests for comparing astrophysical models and observations. MNRAS 524 (1), pp. 1061–1074. External Links: Document, 2302.09308, ADS entry Cited by: Appendix C, §4.2.
- Predicting gravitational wave signals from BPASS white dwarf binary and black hole binary populations of a Milky Way-like galaxy model for LISA. MNRAS 534 (3), pp. 1707–1728. External Links: Document, 2405.20484, ADS entry Cited by: §1.
- Formation of millisecond pulsars. I. Evolution of low-mass X-ray binaries with P_orb¿ 2 days. A&A 350, pp. 928–944. External Links: Document, astro-ph/9909147, ADS entry Cited by: Figure 8, Appendix A, Appendix A, §2.1, §5.
- Coping with loss. Stability of mass transfer from post-main-sequence donor stars. A&A 669, pp. A45. External Links: Document, 2209.12707, ADS entry Cited by: §1.
- Supernova Type Ia progenitors from merging double white dwarfs. Using a new population synthesis model. A&A 546, pp. A70. External Links: Document, 1208.6446, ADS entry Cited by: footnote 1.
- On the Origin of the Close Pair of Degenerate Dwarfs L:870-2. Soviet Astronomy Letters 14, pp. 265. External Links: ADS entry Cited by: §1.
- Testing eccentricity pumping mechanisms to model eccentric long-period sdB binaries with MESA. A&A 579, pp. A49. External Links: Document, 1505.03293, ADS entry Cited by: §5.
@@
- Common Envelope Evolution Redux. In Astrophysics and Space Science Library, E. F. Milone, D. A. Leahy, and D. W. Hobill (Eds.), Astrophysics and Space Science Library, Vol. 352, pp. 233. External Links: Document, 0704.0280, ADS entry Cited by: §1.
- Hot Degenerates in the Montreal-Cambridge-Tololo Survey. II. Two New Hybrid White Dwarfs, MCT 0128-3846 and MCT 0453-2933, and the Nature of the DAB Stars. ApJ 429, pp. 369. External Links: Document, ADS entry Cited by: Table 1.
- On the Formation of Double White Dwarfs through Stable Mass Transfer and a Common Envelope. ApJ 744 (1), pp. 12. External Links: Document, 1102.1039, ADS entry Cited by: §1.
- The gravitational wave signal from diverse populations of double white dwarf binaries in the Galaxy. A&A 521, pp. A85. External Links: Document, 1007.4267, ADS entry Cited by: §1.
- WD mass and orbital period relation of sdB + He WD binaries. MNRAS 505 (3), pp. 3514–3519. External Links: Document, 2106.04361, ADS entry Cited by: §2.1, §5.
- Constraints on Common Envelope Ejection from Double Helium White Dwarfs. ApJ 977 (1), pp. 24. External Links: Document, 2410.14151, ADS entry Cited by: §1, §2.1, §2.1, Figure 3, §4.2, §5, §5.
@@ Appendix A Period/radius - core mass relations
Several studies, in particular into the formation of millisecond radio pulsars with WD companions have determined the period – core mass/WD mass relation that is expected after a phase of stable mass transfer. Rappaport et al. (1995) gave a simple relation, but Tauris and Savonije (1999) and Lin et al. (2011) found that at low mass that relation overestimates the periods. In addition, the radius of the giants and thus the period – core mass relation depends on metallicity. Rappaport et al. (1995) and Tauris and Savonije (1999) find that population II stars, with assumed metallicity of 0.001, i.e about 1/20 of Solar, have about 40% smaller radii. However, such low metallicity is very rare in the Milky Way. In Fig. 7 we plot the cumulative distribution of the metallicity of the total amount of star formation in the Milky Way, based on the derived star formation rate and metallicity by Snaith et al. (2015). It can be seen that there is a similar amount of star formation with metallicities higher than Solar as there is with metallicities below , which on a logarithmic scale lies between the two metallicities used in the models by Tauris and Savonije (1999). We therefore use two metallicities in this study: Solar and intermediate in which the radii are 20% smaller than the Solar values. The latter is similar to ”Pop I+II” in Tauris and Savonije (1999), For the population studies we also use these two metallicities as metallicity bins. The lower metallicity covers about 1/3 of the total SFR (see Fig. 7).
We adapted the curve of Lin et al. (2011) that is given only for solar metallicity and for relatively low mass to match the ones of Rappaport et al. (1995); Tauris and Savonije (1999) at higher mass. For this we multiply their relation with a factor 1.1 for solar metallicity and a factor 0.8 for intermediate metallicity. We note that Gao and Li (2023) find somewhat larger periods.